OpenStudy (anonymous):

Can you please help me with this? the selling price of an item is $100.00 and a profit of 25% was made what was the cost of the item

OpenStudy (kropot72):

The selling price is 100 + 25 = 125% of the cost price. One percent of the cost price is 100/125. 100% of the cost price is \[\frac{100}{125}\times 100=you\ can\ calculate\]

OpenStudy (kropot72):

@Gweng Can you calculate the cost of the item?

OpenStudy (anonymous):

is it 80.00

OpenStudy (kropot72):

Good work! The cost of the item is $80.00.

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

can you do factorization for i have three question on it

OpenStudy (kropot72):

Sorry, I must log out now. Please post your other questions and other people can help.

OpenStudy (anonymous):

ok thanks

OpenStudy (kropot72):

You're welcome :)